OpenStudy (anonymous):

Find the surface area of a rectangular prism that is 16 inches long , 12 inches wide, and 5 inches high

OpenStudy (anonymous):

S.A = 2(lw + lh + wh) S.A = 2((16 x 12) + (16 x 5) + (12 x 5) S.A = 2(192 + 80 + 60) S.A. = 2(332) S.A. = 664\[inches^{2}\]

OpenStudy (anonymous):

664 square inches...